# U-PINet: Physics-Informed Hierarchical Learning for Accurate and Fast 3D RCS Prediction

### v2

@@ Abstract
Accurate radar cross section (RCS) computation is a fundamental task in radar engineering and electromagnetic (EM) scattering analysis, underpinning target signature characterization, detection, and recognition. Conventional computational electromagnetics (CEM) solvers provide high-fidelity RCS predictions but suffer from prohibitive computational costs when applied to $3$-dimensional (3D) targets under multi-aspect configurations. In contrast, purely data-driven neural networks offer high efficiency yet often lack physical consistency and generalization capability. To address these challenges, this paper proposes a U-shaped Physics-Informed Network (U-PINet). To the best of our knowledge, it is the first framework to establish a fully end-to-end, physics-informed hierarchical architecture for fast and accurate RCS computation, grounded in the governing principles of CEM. Inspired by the near–far field decomposition in classical fast solvers, U-PINet explicitly models local EM coupling and long-range radiation effects through a hierarchical operator design. A physics-guided graph construction is further introduced to represent self- and mutual-coupling among mesh elements of complex 3D targets, enabling physically interpretable intermediate representations. By embedding EM governing equations as residual constraints, the proposed framework achieves end-to-end, physically consistent RCS prediction with significantly improved computational efficiency. Extensive numerical experiments demonstrate that U-PINet attains solver-level RCS accuracy with orders-of-magnitude runtime reduction, while exhibiting strong generalization to unseen target geometries under limited training data.
@@ I Introduction
Radar cross section (RCS) computation plays a fundamental role in radar engineering and electromagnetic (EM) analysis, serving as a primary descriptor of target scattering characteristics for detection [1], recognition [2], and signature evaluation [3]. Accurate RCS prediction establishes the quantitative relationship between the incident EM waves and the physical properties of scattering targets [4]. To achieve physically consistent RCS modeling, EM scattering is typically formulated as a boundary-value problem governed by Maxwell’s equations.
The conventional approach to RCS computation relies on numerically solving Maxwell’s equations to determine the induced surface currents on scattering objects, from which far-field scattering responses and RCS values are subsequently evaluated. Classical computational electromagnetics (CEM) methods, such as the Method of Moments (MoM) [5], Finite Element Method (FEM) [6], and Finite-Difference Time-Domain (FDTD) [7], have been extensively investigated and demonstrated high accuracy in modeling EM scattering phenomena. Among these techniques, MoM has been particularly validated for its accuracy in handling targets with complex geometries [8], making it a widely adopted tool for high-fidelity RCS prediction.
@@
Within the conventional MoM framework, extensive efforts have been devoted to improving the efficiency of large-scale RCS computation. Structural algorithmic enhancements [9, 10] and the utilization of high-performance computing platforms [11, 12] have been proposed to alleviate the computational burden associated with dense matrix operations. Among these advances, the Multilevel Fast Multipole Algorithm (MLFMA) [13] introduces a hierarchical decomposition with near- and far-field partitioning to significantly reduce the computational complexity of MoM-based solvers. Further developments in kernel compression and fast matrix–vector multiplication [14, 15, 16] have enabled more efficient evaluation of far-field interactions and reduced memory consumption.
Despite these notable improvements, classical solvers remain challenged by their high computational cost and limited flexibility, particularly in RCS analysis involving large-scale three-dimensional targets, multi-aspect scanning, or repeated evaluations. Such scenarios are common in practical radar applications, yet the deterministic structure and matrix-based formulation of conventional solvers fundamentally constrain their scalability and efficiency.
@@
In recent years, deep learning techniques have been explored as an alternative paradigm for accelerating RCS computation and EM scattering prediction. By leveraging their strong nonlinear approximation capability, neural networks can learn complex scattering behaviors from data and provide rapid inference once trained. Various end-to-end learning frameworks have been proposed to infer EM responses directly from geometric representations, including supervised descent approaches mapping image-based target descriptions to scattering characteristics [17], multi-task learning architectures that enhance generalization via shared latent spaces [18], and models incorporating extended input modalities such as radiation-direction maps [19]. Moreover, specific network architectures have been tailored for efficient field solving: fully convolutional networks (ConvNet) have been applied to predict potential distributions in 3D models with high numerical accuracy [20], while physics-augmented U-Net structures have been utilized to ensure physical consistency in EM field prediction [21].
However, most existing learning-based approaches for RCS and EM scattering prediction operate without explicit physical constraints. As a consequence, they function as black-box predictors whose internal representations lack correspondence to underlying electromagnetic principles. This deficiency raises concerns regarding their reliability and generalization, especially for targets with complex geometries or under unseen electromagnetic configurations [22]. Therefore, despite their promising efficiency, purely data-driven models remain insufficient for physically sensitive RCS computation tasks.
@@
To overcome the computational and generalization challenges in RCS computation, Physics-Informed Neural Networks (PINNs) [23] have been proposed to embed Maxwell’s equations and boundary conditions directly into the learning framework, thereby constraining the solution space to physically consistent regimes. Typically, PINNs enforce physical laws by incorporating governing equations into the loss function or by coupling neural networks with physics-based operators, such that the learned representations adhere to EM principles throughout training and inference. From an RCS perspective, this physics-informed paradigm is especially beneficial, as it not only stabilizes intermediate field or current predictions but also enhances the fidelity and robustness of the resulting far-field scattering signatures, enabling accurate RCS curve prediction with reduced training data.
@@
Other efforts extend physics-informed learning to multilevel fast algorithms. Under the MLFMA framework, the far-field translation operator has been approximated by radial basis function networks to capture long-range interactions between distant clusters. Building on this idea, [33] proposes a hybrid deep model that jointly trains multilevel translation operators through an embedding layer and a self-attention mechanism, enabling the automatic learning of dependencies across hierarchical levels [34]. The learning-assisted framework in [35] further accelerates MLFMA translation operations by modeling EM interactions in a coarse-to-fine manner, reducing computational complexity for large-scale scattering problems. These developments indicate that physics-guided learning can effectively accelerate the evaluation of long-range radiation effects, which are directly related to the formation of multi-aspect far-field signatures and thus the resulting RCS curves.
Overall, existing PINN-based approaches can accelerate EM simulations by embedding physical priors, but they are mostly implemented as auxiliary modules within conventional solver pipelines. Consequently, mesh generation, matrix assembly, and iterative residual updates remain required, and expensive re-computation is often unavoidable when geometries or illumination parameters change. This limitation is particularly problematic for RCS analysis under multi-aspect, where RCS curves must be evaluated repeatedly and efficiently. Hence, the scalability and generalization of such hybrid schemes are still insufficient for large-scale or time-sensitive RCS computation.
Inspired by these recent advances, we propose a U-shaped Physics-Informed Network (U-PINet), the first fully deep learning-based, physics-informed hierarchical framework that integrates electromagnetic physical priors into every stage of the network for RCS computation. This enables end-to-end modeling that replaces conventional solvers while preserving physical consistency and ensuring scalability. The main technical contributions of the paper are summarized as follows:
  • A MoM-guided sparse Graph Neural Network (GNN) enabling efficient EM-based computation is devised. The proposed incorporation of MoM schemes allows the considered GNN to simulate local self/mutual coupling among its mesh elements in a physically interpretable and computationally efficient manner, thus supporting accurate RCS prediction.
  • By introducing multiscale encoding and skipping connections between near-/far field modules, a novel hierarchical near-/far-field fusion mechanism that preserves both global consistency and local details is presented, enabling stable far-field scattering responses and RCS signatures.
  • Our numerical investigations showcase that the proposed U-PINet achieves solver-level accuracy on RCS prediction with significantly lower computational cost, while offering satisfactory generalization across diverse $3$-Dimensional (3D) scattering scenarios, thus enabling efficient and scalable RCS curve modeling.
Figure 1: Illustration of the RCS calculation for a representative target.
@@
Figure 2: Framework Comparison for RCS Prediction. The diagram illustrates the pathway to the final RCS calculation ($\sigma$) using the standard MLFMA approach (top) versus the proposed U-PINet method (bottom). U-PINet utilizes physics-informed learning to directly resolve field interactions, bypassing traditional iterative solvers to rapidly determine the scattered field for RCS assessment.
@@
The remainder of the paper is organized as follows. Section II reviews the MLFMA foundations and introduces the proposed U-PINet principle discussing its core feature of preserving physical structure while avoiding solver-dependent bottlenecks. Section III outlines the proposed physics-guided neural network architecture for 3D EM scattering modeling, in particular, its near-field graph module and the hierarchical far-field learning mechanism. Section IV introduces our experimental setup and Section V presents a series of experiments assessing our U-PINet’s modeling performance. Section VI concludes the paper and highlights future research directions.
II Physics-Guided Design of U-PINet
II-A Problem Formulation and RCS Definition
@@
Consider a representative perfectly electrically conducting (PEC) target with surface $\mathcal{D}$ situated in a homogeneous background medium, as illustrated in Fig. 1. Without loss of generality, we construct a three-dimensional (3D) coordinate system where the target is centered at the origin $\boldsymbol{o}$. Let an observation point be located at $\boldsymbol{s}$ with distance $s$, elevation angle $\theta$, and azimuth angle $\phi$, which can be parameterized as $\boldsymbol{s}=s(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$.
@@
Let $\boldsymbol{p},\boldsymbol{p}^{\prime}\in\mathcal{D}$ denote an observation point and a source point on the target surface, respectively. According to electromagnetic theory, the tangential component of the total electric field must vanish on the PEC surface $\mathcal{D}$. This boundary condition is expressed as:
$\hat{\boldsymbol{n}}(\boldsymbol{p})\times\left(\boldsymbol{E}^{\mathrm{inc}}(\boldsymbol{p})+\boldsymbol{E}^{\mathrm{sca}}(\boldsymbol{p})\right)=\boldsymbol{0},\qquad\forall\boldsymbol{p}\in\mathcal{D},$ (1)
@@
where $\hat{\boldsymbol{n}}(\boldsymbol{p})$ is the outward unit normal vector. Enforcing (1) together with the radiation condition yields the Electric Field Integral Equation (EFIE). Specifically, under a prescribed incident field $\boldsymbol{E}^{\mathrm{inc}}(\boldsymbol{p})$, the induced surface current density $\boldsymbol{J}(\cdot)$ satisfies:
$\int_{\boldsymbol{p}^{\prime}\in\mathcal{D}}\mathbf{G}(\boldsymbol{p},\boldsymbol{p}^{\prime})\,\boldsymbol{J}(\boldsymbol{p}^{\prime})\,\mathrm{d}\boldsymbol{p}^{\prime}=\boldsymbol{E}^{\mathrm{inc}}(\boldsymbol{p}),\qquad\forall\boldsymbol{p}\in\mathcal{D},$ (2)
@@
where $\mathbf{G}(\boldsymbol{p},\boldsymbol{p}^{\prime})$ denotes the dyadic Green’s function kernel in the background medium [36]. The induced surface current density $\boldsymbol{J}(\cdot)$ is the unknown quantity to be solved for in (2).
@@
Once the induced current is determined, it radiates the scattered field to the exterior region. The scattered electric field at the observation point $\boldsymbol{s}$ is given by the radiation integral:
@@
Under far-field conditions, the observation point $\boldsymbol{s}$ is sufficiently distant from the target. Consequently, the distance and direction from $\boldsymbol{s}$ to any point $\boldsymbol{p}$ on the target can be approximated as $\|\boldsymbol{p}-\boldsymbol{s}\|_{2}\to\|\boldsymbol{o}-\boldsymbol{s}\|_{2}=\|\boldsymbol{s}\|_{2}$ and $(\theta,\phi)$, respectively. Based on this approximation, the bistatic Radar Cross Section (RCS) is mathematically defined as [4]:
$\sigma(\theta,\phi)=\lim_{s\rightarrow\infty}\left(4\pi s^{2}\cdot\left|\frac{\boldsymbol{E}^{\mathrm{sca}}(\boldsymbol{s})}{\boldsymbol{E}^{\mathrm{inc}}(\boldsymbol{o})}\right|^{2}\right),$ (4)
where $\boldsymbol{E}^{\mathrm{inc}}(\boldsymbol{o})$} is the amplitude of the plane wave incident from a fixed direction $(\theta_{\mathrm{inc}},\phi_{\mathrm{inc}})$. Note that for monostatic RCS, the observation angle $(\theta,\phi)$ coincides with the backscattering direction of the incident wave.
II-B Methodological Foundation Inspired by MLFMA
To motivate the methodological structure underlying our model design, we briefly review the physics-based computational paradigm inspired by MLFMA. As indicated by (2)–(4), accurate evaluation of the RCS signature $\sigma(\theta,\phi)$ requires the induced surface current $\boldsymbol{J}$, which is obtained by solving the EFIE in (2).
To this end, we discretize (2) using the Rao–Wilton–Glisson (RWG) basis functions defined on mesh edges [37]. The resulting MoM formulation yields a dense linear system $\mathbf{Z}\boldsymbol{I}=\boldsymbol{V}$, whose direct matrix–vector products scale as $\mathcal{O}(N^{2})$, with $N$ denoting the number of unknown degrees of freedom, i.e., the number of RWG basis functions (typically on the order of the number of interior mesh edges). Here, $\mathbf{Z}\in\mathbb{C}^{N\times N}$ is the impedance matrix, $\boldsymbol{I}\in\mathbb{C}^{N}$ collects the RWG expansion coefficients of the induced surface current, and $\boldsymbol{V}\in\mathbb{C}^{N}$ contains the tested incident field. To enable fast evaluation in the MLFMA, the impedance matrix is decomposed as follows:

### v4

# U-PINet: Physics-Informed Hierarchical Learning for Radar Cross Section Prediction via 3D Electromagnetic Scattering Reconstruction

@@ Abstract
Conventional computational electromagnetics (CEM) solvers can deliver high fidelity radar cross section (RCS) signatures by first solving the induced surface currents on $3$-dimensional (3D) targets and then evaluating the scattered fields via radiation integrals. However, their computational cost becomes prohibitive for repeated queries and large-scale 3D scenarios. Recent purely data-driven networks improve efficiency, yet they often bypass this scattering mechanism, which may compromise physical consistency and generalization. To bridge this gap, in this paper, we propose U-PINet, a fully end-to-end, physics-informed hierarchical network for efficient RCS prediction via 3D electromagnetic scattering reconstruction. Once the scattering quantities are reconstructed, scattered fields and RCS can be evaluated for arbitrary observation directions via the radiation integral. U-PINet explicitly learns physics-consistent intermediate scattering representations by modeling local electromagnetic coupling and long-range radiation effects through a hierarchical operator design inspired by near–far field decomposition in fast solvers. A physics-guided graph neural network is incorporated to capture self- and mutual-coupling among mesh elements of complex targets, enabling physically interpretable intermediate representations. By embedding governing equations as residual constraints, U-PINet enables accurate object reconstruction of scattering quantities and consequently reliable RCS prediction across observation directions, while significantly reducing runtime. Extensive numerical experiments demonstrate that U-PINet achieves EM-solver-level RCS accuracy and 3D object reconstruction with orders of magnitude speedups, and generalizes well to unseen geometries under limited training data.
@@ I Introduction
Radar cross section (RCS) computation plays a fundamental role in radar engineering and electromagnetic (EM) analysis, serving as a primary descriptor of target scattering characteristics for detection [1], recognition [2], and signature evaluation [3]. Accurate RCS prediction establishes the quantitative relationship between the incident EM waves and the physical properties of scattering targets [4]. From a physics standpoint, RCS constitutes an engineering-level observable that follows from the underlying EM scattering quantities, most notably the induced surface currents governed by Maxwell’s equations [5]. Once these induced currents are obtained, the scattered field can be evaluated for arbitrary observation directions via radiation integrals, from which the corresponding RCS is deduced. Therefore, accurate $3$-dimensional (3D) EM scattering reconstruction of the induced currents is key enabler for fast and high fidelity RCS prediction.
@@
Classical computational electromagnetics (CEM) methods, such as the Method of Moments (MoM) [6], Finite Element Method (FEM) [7], and Finite-Difference Time-Domain (FDTD) [8], have been extensively investigated, demonstrating high accuracy in modeling EM scattering phenomena. Among these techniques, MoM has been particularly efficient for handling targets with complex geometries [9], providing increased accuracy, making it a widely adopted tool for high fidelity EM scattering reconstruction and subsequent RCS evaluation. In addition, to alleviate the computational burden associated with dense matrix operations in large-scale RCS computation cases, structural algorithmic enhancements of MoM [10, 11] in conjunction with the utilization of high-performance computing platforms [12, 13] have been proposed. The Multilevel Fast Multipole Algorithm (MLFMA) [14] introduces a hierarchical decomposition with near-/far-field partitioning to significantly reduce the computational complexity of MoM-based solvers, thereby accelerating the evaluation of long-range radiation effects that underpin scattered fields and RCS signatures. Further developments in kernel compression and fast matrix–vector multiplication [15, 16, 17] have enabled more efficient evaluation of far-field interactions and reduced memory consumption. However, despite these notable improvements, classical solvers remain challenged by their high computational cost and limited flexibility, and particularly in RCS analysis involving large-scale 3D targets, multi-aspect scanning, or repeated evaluations. Such scenarios are common in practical radar applications, yet the deterministic structure and matrix-based formulation of conventional solvers fundamentally constrain their scalability and efficiency.
@@
In recent years, deep learning techniques have been explored as an alternative paradigm for accelerating EM scattering prediction and RCS computation. By leveraging their strong nonlinear approximation capability, neural networks can learn complex scattering behaviors from data and provide rapid inference once trained. Various end-to-end learning frameworks have been proposed to infer EM responses directly from geometric representations, including supervised descent approaches mapping image-based target descriptions to scattering characteristics [18], multi-task learning architectures that enhance generalization via shared latent spaces [19], and models incorporating extended input modalities such as radiation-direction maps [20]. Moreover, specific network architectures have been tailored for efficient field solving: fully convolutional networks (ConvNet) have been applied to predict potential distributions in 3D models with high numerical accuracy [21], while physics-augmented U-Net structures have been utilized to ensure physical consistency in EM field prediction [22]. However, most existing learning-based approaches for RCS and EM scattering prediction operate without explicit physical constraints. As a consequence, they function as black-box predictors whose internal representations lack correspondence to underlying EM principles. This deficiency raises concerns regarding their reliability and generalization, especially for targets with complex geometries or under unseen EM configurations [23]. Therefore, despite their promising efficiency, purely data-driven models remain insufficient for physically sensitive RCS computation tasks.
@@
To overcome the computational and generalization challenges in EM scattering reconstruction, Physics-Informed Neural Networks (PINNs) [24] have been proposed to embed Maxwell’s equations and boundary conditions directly into the learning framework, thereby constraining the solution space to physically consistent regimes. Typically, PINNs enforce physical laws by incorporating governing equations into the loss function or by coupling neural networks with physics-based operators, such that the learned representations adhere to EM principles throughout training and inference. From an RCS perspective, this physics-informed paradigm is especially beneficial, as it not only stabilizes intermediate field or current predictions but also enhances the fidelity and robustness of the resulting far-field scattering signatures, enabling accurate RCS curve prediction with reduced training data.
@@
Other efforts extend physics-informed learning to multilevel fast algorithms. Under the MLFMA framework, the far-field translation operator has been approximated by radial basis function networks to capture long-range interactions between distant clusters. Building on this idea, [34] proposes a hybrid deep model that jointly trains multilevel translation operators through an embedding layer and a self-attention mechanism, enabling the automatic learning of dependencies across hierarchical levels [35]. The learning-assisted framework in [36] further accelerates MLFMA translation operations by learning EM interactions in a coarse-to-fine manner, reducing computational complexity for large-scale scattering problems. These developments indicate that physics-guided learning can effectively accelerate the evaluation of long-range radiation effects, which are directly related to the formation of multi-aspect far-field signatures and thus the resulting RCS curves.
Overall, existing PINN-based EM solvers inject physical priors, but remain tethered to conventional pipelines. This implies that mesh generation, matrix assembly, and iterative residual updates still dominate cost, making multi-aspect RCS calculation and repeated evaluations particularly inefficient. Motivated by this fact, in this paper, we propose U-PINet, a fully end-to-end physics-informed hierarchical framework that reconstructs the induced surface currents and then obtains the scattered fields and RCS signatures via the governing radiation formulations for arbitrary observation directions, thereby supporting diverse sensing geometries, while replacing classical solvers with physically consistent and scalable inference. The main technical contributions are summarized as follows.
  • A physics-guided graph neural network (GNN) is developed, which serves as a near-field coupling engine efficiently capturing self- and mutual-interactions among mesh elements and producing physically interpretable local current representations.
  • A hierarchical near-/far-field fusion operator with multiscale encoding and skip connections is introduced to integrate the learned near-field couplings with long-range radiation effects. This operator enables reconstructing the induced surface currents, yielding stable scattered field and RCS evaluation via the radiation integral across arbitrary observation directions.
  • Our extensive numerical experiments demonstrate that U-PINet achieves EM-solver-level accuracy for both induced-current reconstruction and RCS prediction with orders of magnitude runtime reduction, offering also strong generalization.
Figure 1: Geometry for EM scattering and RCS evaluation. A PEC target with surface $\mathcal{D}$ (point $\boldsymbol{p}$ belongs in this surface) is illuminated by an incident plane wave $\boldsymbol{E}^{\mathrm{inc}}$ from direction $(\theta_{\mathrm{inc}},\phi_{\mathrm{inc}})$, and the scattered field $\boldsymbol{E}^{\mathrm{sca}}$ is observed at a far-field point $\boldsymbol{s}$ along the direction $(\theta,\phi)$.
@@
Figure 2: The RCS, $\sigma(\theta,\phi)$, calculation procedure using the standard MLFMA framework [14] (top) and the proposed U-PINet (bottom), which utilizes physics-informed learning to directly resolve field interactions. In this way, traditional iterative solvers are bypassed enabling to rapidly determine the scattered field for the desired RCS assessment. U-PINet’s core components and overall architecture are detailed in Sections III and IV.
@@
The remainder of the paper is organized as follows. Section II reviews the MLFMA foundations and introduces the proposed U-PINet principle discussing its core feature of preserving physical structure while avoiding solver-dependent bottlenecks. Section IV outlines the proposed physics-guided neural network architecture for 3D EM scattering reconstruction, in particular, its near-field graph module and the hierarchical far-field learning mechanism. Section V introduces our experimental setup and Section VI presents a series of experiments assessing our U-PINet’s reconstruction performance. Section VII concludes the paper and highlights future research directions.
II Problem Formulation and State of the Art
II-A Scattered Electric Field and Radar Cross Section
@@
Consider a representative perfectly electrically conducting (PEC) target with surface $\mathcal{D}$ embedded in a homogeneous background medium, as illustrated in Fig. 1. Without loss of generality, we adopt a 3D Cartesian coordinate system with the target centered at the origin $\boldsymbol{o}$. The target is illuminated by a prescribed incident plane wave $\boldsymbol{E}^{\mathrm{inc}}$ impinging from direction $(\theta_{\mathrm{inc}},\phi_{\mathrm{inc}})$, and the scattered response is observed at an exterior far-field point $\boldsymbol{s}$ parameterized by range–angle coordinates as $\boldsymbol{s}=s(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$. Under far-field conditions ($s\rightarrow\infty$), $\boldsymbol{s}-\boldsymbol{p}||_{2}\approx s$ holds for all $\boldsymbol{p}\in\mathcal{D}$, so the observation is fully characterized by the angles $(\theta,\phi)$.
@@
Let $\boldsymbol{p},\boldsymbol{p}^{\prime}\in\mathcal{D}$ denote a surface testing point and a surface source (integration) point on the target, respectively. According to EM theory, the tangential component of the total electric field must vanish on the PEC surface $\mathcal{D}$. Using $\hat{\boldsymbol{n}}(\boldsymbol{p})$ to denote the outward unit normal vector at the testing point $\boldsymbol{p}\in\mathcal{D}$, this boundary condition is mathematically expressed as follows:
$\hat{\boldsymbol{n}}(\boldsymbol{p})\times\left(\boldsymbol{E}^{\mathrm{inc}}(\boldsymbol{p})+\boldsymbol{E}^{\mathrm{sca}}(\boldsymbol{p})\right)=\boldsymbol{0}.$ (1)
@@
Enforcing this condition together with the radiation condition yields the EFIE. Specifically, under a prescribed incident field $\boldsymbol{E}^{\mathrm{inc}}(\boldsymbol{p})$, the induced surface current density $\boldsymbol{J}(\cdot)$ satisfies the following expression $\forall\boldsymbol{p}\in\mathcal{D}$:
$\int_{\boldsymbol{p}^{\prime}\in\mathcal{D}}\mathbf{G}(\boldsymbol{p},\boldsymbol{p}^{\prime})\,\boldsymbol{J}(\boldsymbol{p}^{\prime})\,\mathrm{d}\boldsymbol{p}^{\prime}=\boldsymbol{E}^{\mathrm{inc}}(\boldsymbol{p}),$ (2)
@@
where $\mathbf{G}(\cdot,\cdot)$ denotes the dyadic Green’s function kernel in the background medium [37]. The induced surface current density $\boldsymbol{J}(\cdot)$ is the unknown quantity to be solved for in (2).
@@
Once the induced surface current density $\boldsymbol{J}$ is obtained, the scattered field can be evaluated at an arbitrary exterior observation point $\boldsymbol{s}$ via the radiation integral:
@@
The corresponding RCS at $(\theta,\phi)$ is then defined as [4]:
$\sigma(\theta,\phi)=\lim_{s\rightarrow\infty}\left(4\pi s^{2}\left|\frac{\boldsymbol{E}^{\mathrm{sca}}(\boldsymbol{s})}{\boldsymbol{E}^{\mathrm{inc}}(\boldsymbol{o})}\right|^{2}\right),$ (4)
where $\boldsymbol{E}^{\mathrm{inc}}(\boldsymbol{o})$} denotes the amplitude of the incident plane wave at the origin. It is noted that different sensing setups are implied through different combinations of the illumination and observation directions. Monostatic RCS corresponds to the backscattering direction of $(\theta_{\mathrm{inc}},\phi_{\mathrm{inc}})$, whereas bistatic/multi-static configurations simply correspond to other receive directions. Therefore, once the induced surface current $\boldsymbol{J}$ is accurately reconstructed, the RCS can be evaluated for arbitrary observation angles via the radiation integral.
II-B The MLFMA Framework
To motivate the methodological structure underlying our model design, we briefly overview the physics-based computational paradigm of MLFMA [14]. As indicated by (2)–(4), accurate evaluation of the RCS signature $\sigma(\theta,\phi)$ requires the induced surface current $\boldsymbol{J}$, which is obtained by solving the EFIE in (2). To this end, expression (2) can be discretized using the Rao–Wilton–Glisson (RWG) basis functions defined on mesh edges [38]. The resulting MoM formulation yields a dense linear system $\mathbf{Z}\boldsymbol{I}=\boldsymbol{V}$, whose direct matrix–vector products scale as $\mathcal{O}(N^{2})$, with $N$ denoting the number of unknown degrees of freedom, i.e., the number of RWG basis functions (typically on the order of the number of interior mesh edges). In the latter system, $\mathbf{Z}\in\mathbb{C}^{N\times N}$ is the impedance matrix, $\boldsymbol{I}\in\mathbb{C}^{N}$ collects the RWG expansion coefficients of the induced surface current, and $\boldsymbol{V}\in\mathbb{C}^{N}$ contains the tested incident field. To enable fast evaluation in the MLFMA, the impedance matrix is decomposed as follows: